Grade 12th passPhysical Chemistry

Show that the velocity of an electron in the first bohrs orbit is (1/137)c where c is velocity of light

Sambhaji patil
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
let us use the Bohr's hypothesis 
Angular momentum of electron is integral multiple of h/2pi 
So m v r = h / 2 pi 
Now m = 9.1 * 10^-31 kg 
r = 0.54 * 10^-10 m 
h = 6.646 * 10^-32 J s 
and pi = 3.14 
Now you can compute v 
Multiply by c and divide it by 3 * 10^8  
You would get the required one and that will be 1/137 * C
 
Regards
Arun